The real line with the usual topology is locally compact
1
Expert's answer
2021-09-13T00:04:50-0400

"\text{$\mathbb{R}$ is locally compact since given $\epsilon > 0$ and x $\in \mathbb{R},$ the neighbourhood $(x-\epsilon, x+\epsilon )$}\\
\text{is contained in $[x-\epsilon, x+\epsilon ]$ and it is known that $[x-\epsilon, x+\epsilon ] \subset \mathbb{R}$ is compact. }\\
 \text{Hence showing that $\mathbb{R}$ is locally compact.}"